# SG-Reg: Generalizable and Efficient Scene Graph Registration

### v1

@@ VII-G Limitations
Another limitation of the proposed method is that it cannot register highly cluttered scenes, such as real-world industrial environments. The reconstructed semantic scene graphs in these scenarios contain significant semantic noise. Due to the large sizes of these scenes, directly matching their semantic objects is challenging. One potential direction for improvement is to further partition an industrial scene into smaller areas, such as rooms. This would allow us to search for correspondences from the top layers down to the bottom layers. We will explore solutions to address these limitations in our future works.
@@ VIII Conclusion
In this work, we propose a learning-based registration method that aligns two semantic scene graphs without requiring an initial value. During the neural network stage, the model encodes multiple modalities of a semantic node, including its semantic label, local topology with spatial awareness, and geometric shape. It also incorporates the dense point cloud associated with each semantic node. The network then searches for correspondences at two levels: coarse semantic nodes and dense point clouds. In the optimization stage, we employ a robust pose estimator to compute the transformation based on these correspondences. Thanks to automatic data generation, we train the network in a self-supervised manner without the need for any ground-truth semantic annotations. In the evaluation, we compare our method with point cloud registration networks and semantic registration networks. Our approach achieves a higher recall rate, particularly in medium and large scenes. Additionally, our GPU inference requires less than $2.2$ GFLOPS, while the baseline can require up to $1,314$ GFLOPS. We deploy SG-Reg in a two-agent SLAM system to align the coordinates between the two agents. In real-world indoor scenes, SG-Reg significantly outperforms handcrafted descriptors in Hydra. When compared to the state-of-the-art image matching network HLoc, SG-Reg achieves a similar registration success rate. However, in frames with large viewpoint differences, SG-Reg registers the scene more accurately and robustly than HLoc. Additionally, our communication module requires only 3.9% of the bandwidth that HLoc demands. On the other hand, we recognize the system demands heavy front-end modules to reconstruct the semantic maps. In the two-agent benchmark, we ignore the pose drift within each agent and register their maps as two rigid bodies. Despite these limitations, our system is generalizable, offers higher computational efficiency during inference, and communicates data using lower bandwidth.

### v2

@@ VII-G Limitations
Moreover, we find some failed registration in the long-tailed scenes, such as empty rooms and scenes in symmetric layout. We provide their results in the supplementary material, which can be further improved in the future.
@@ VIII Conclusion
In this work, we propose a learning-based registration method that aligns two semantic scene graphs without requiring an initial value. The neural network encodes coarse semantic node features, which are fused from their semantic label, local topology with spatial awareness, and geometric shapes. Besides, our network encodes the dense point cloud that are belong to each semantic node. Then, the network searches for correspondences at two levels: coarse semantic nodes and dense point clouds. In the optimization stage, we employ a robust pose estimator to compute the transformation based on these correspondences. Thanks to automatic data generation, we train the network in a self-supervised manner without the need for any ground-truth semantic annotations. In the evaluation, we compare our method with point cloud registration networks and semantic registration networks. Our approach achieves a higher recall rate, particularly in medium and large scenes. Additionally, our GPU inference requires less than $2.2$ GFLOPS, while the baseline can require up to $1,314$ GFLOPS. We deploy SG-Reg in a two-agent SLAM system to align the coordinates between the two agents. In real-world indoor scenes, SG-Reg significantly outperforms handcrafted descriptors in Hydra. When compared to the state-of-the-art image matching network HLoc, SG-Reg achieves a similar registration success rate. However, in frames with large viewpoint differences, SG-Reg registers the scene more accurately and robustly than HLoc. Additionally, our communication module requires only 3.9% of the bandwidth that HLoc demands. On the other hand, we recognize the system demands heavy front-end modules to reconstruct the semantic maps. In the two-agent benchmark, we ignore the pose drift within each agent and register their maps as two rigid bodies. Despite these limitations, our system is generalizable, offers higher computational efficiency during inference, and communicates data using lower bandwidth.
Appendix A Scene Graph Network
A-1 Multi-stage Training
We utilize a multi-stage training strategy to reduce data complexity and improve training efficiency. Firstly, BERT[24] is pre-trained and fixed. Secondly, we pre-train the shape encoder in the ScanNet dataset, ensuring the point cloud backbone (KPConv) and the object shape backbone effectively capture geometric features. The shape loss in Equation (16) supervises the pre-training. Thirdly, we train all of the network layers end-to-end, including the triplet-GNN and the hierarchical matching layers. Since the scene graph network encodes multiple modalities, pre-training the features from semantic labels and point cloud benefits the end-to-end training.
A-2 Ablation study
Num. of Correspondences IR(%) PIR(%) RR(%)
$K_{p}=256$ 492 19.6 51.6 78.0
$K_{p}=512$ 313 26.8 55.0 69.0
$K_{p}=1024$ 126 36.9 55.9 59.0
TABLE XVII: Ablation study on sampling points $K_{p}$ from each semantic node. We evaluate the performance in cross-domain benchmark.
As shown in TABLE XVII, we set parameter $K_{p}$ to be a larger value. It is the number of points we sampled from each semantic node, as explained in Sec. IV-C. In a larger value of $K_{p}$, we have three interesting findings. Firstly, the number of correspondences declines. It is because the optimal transport algorithm[26] run multiple normalization operations. A growing size of the similarity matrix suppress the average similarity scores after normalizations. Hence, fewer correspondences are detected. Secondly, the inlier ratio grows higher as it implicitly sets a higher threshold. Thirdly, the final registration recall declines. It implies that the inlier ratio is not the only key factor that decides the registration accuracy. The back-end estimator also needs a larger number of inliers to prune outliers and predict a precise transformation.
Appendix B Hierarchical communication
B-1 Dense matching Vs Coarse-to-fine matching
(a) Dense matching $\sigma=0.02$; $36,182$ MFLOPS. Refer to caption (b) Dense matching $\sigma=0.05$. Refer to caption (c) Coarse-to-fine (Ours); 602 MFLOPS.
Figure 13: Dense match points and coarse-to-fine match points.
We analysis the advantages of coarse-to-fine matching in Sec. VII-E(10). Compared to dense matching, it requires lower computational FLOPS and result in a higher inlier ratio. As shown in Fig. 13(a), we evaluate the dense matching in a scene. Because of the large amount of similar geometric features, dense matching approach has a low inlier ratio and very few inliers. We set a higher matching threshold in Fig. 13(b) and it can find a larger amount of correspondences. But the inliers ratio is still low. On the other hand, coarse-to-fine approach generates the highest inlier ratio. Its computational FLOPS is also lower, as we analyzed in Sec. VII-E(10).
B-2 How dense matching improves registration
We provide a visualization result to further support the evaluation in TABLE XIII.
Low-overlap Refer to caption Refer to caption
High-overlap Refer to caption Refer to caption (a) Coarse message (a) Dense message
Figure 14: The matching results at a low-overlap scene and a high-overlap scene.
As shown in Fig. 14(a), SG-Reg relies on frames containing only coarse messages for registration. In each frame, the pose estimator uses the centers of matched semantic nodes to align the scenes. While coarse messages enable successful registration in some high-overlap scenes, they are likely to fail in low-overlap scenarios due to the limited number of true positive matches, which poses serious challenges for accurate registration. In contrast, as illustrated in Fig. 14(b), dense messages significantly improves SG-Reg’s ability to register low-overlap scenes by providing more measurements utilizing point correspondences. Additionally, dense messages enhance registration accuracy in high-overlap scenes. This further supports our claim related to TABLE XIII, that broadcasting dense messages improve registration accuracy.
Appendix C Efficiency in high inlier ratio
In Sec. V, we mentioned that high inlier ratio requires a longer computational complexity. To clarify it, we run a simple experiment and show its results.
Figure 15: Computational profile of maximum clique-based registration under varying inlier ratios (total correspondences: 1000). Each inlier ratio was tested ten times with results averaged. Graph construction time (green) grows with inlier ratio due to the addition of graph edges, while maximum clique solving time (blue) first increases then decreases as graph density enables aggressive pruning. Registration error (red) remains low at low inlier ratios, validating the robustness of maximum clique pruning even under extreme outlier rates.
Here, we aim to express that in scenarios with a high inlier ratio, pruning outliers via maximum clique algorithm is less efficient compared with directly employing Graduated Non-Convexity (GNC) for outlier removal and pose estimation. The reason we cite the literature [69] is that the maximum clique solver we use in practice originates from this paper. Although, in theory, the maximum clique algorithm has exponential complexity in the worst case, the algorithm we employ is based on branch-and-bound with efficient pruning strategies. As a result, its complexity initially increases linearly with the inlier ratio but decreases after reaching a certain threshold. This is because, at the beginning, a higher inlier ratio leads to more edges in the graph, requiring the algorithm to explore more possibilities. However, once the inlier ratio exceeds a certain level, it provides a tight upper bound. As illustrated in the literature[69], a tight upper bound facilitates a more effective pruning. To validate this, we conducted an experiment on a registration example from our dataset to investigate the relationship between maximum clique search efficiency and the inlier ratio, as shown in Figure 15. As illustrated, the maximum clique search time aligns with our theoretical analysis.
From a practical perspective, we must also consider the time required to construct the graph. As the number of inliers increases, more edges need to be established, which increases the graph construction time, even with parallelization strategies. Additionally, we present the variation in registration error, demonstrating that even at low inlier ratios, the maximum clique-based outlier removal algorithm maintains robust outlier rejection. Given this finding, we adopt a practical approach: we first attempt registration using GNC. If the inlier ratio calculated from the GNC registration result is sufficiently high, we avoid using the slower maximum clique algorithm.
Appendix D Registrations in semantic noise
Colored point cloud Semantic scene graph Matched nodes
Figure 16: Two successful registration with semantic noise. In the semantic scene graph, we highlight a few of the nodes that are over-segmented, falsely labeled and partially reconstructed.
(a) Low overlapped scenes.
(b) Empty room.
(c) Symmetric couch layout.
Figure 17: Failure registration. In the middle column, we highlight a few significant semantic noise. In the right column, we visualize the point correspondences.
D-1 Success cases
We show two indoor scenes and highlight their semantic noise. As shown in Fig. 16, SG-Reg registers them successfully.
D-2 Failed cases
We further analyze a few failure cases of SG-Reg to illustrate our upper limits. Some of the limitations can be further addressed in our future works. Fig. 17 shows three examples of failed registration. We analyze the reasons behind their failures.
  • Significant Semantic Noise: Semantic noise is prevalent in all the scenes we reconstructed; however, the failed scenes exhibit more pronounced noise. As illustrated in the first column of Fig. 17, these scenes contain numerous objects that are over-segmented and inaccurately labeled, which severely disrupts the spatial topological relationships between objects. Furthermore, many of these scenes feature objects belonging to identical semantic categories and positioned close to one another, leading to inaccurate node matches. This, in turn, results in a substantial number of outliers in the point correspondences, ultimately causing the registration failures.
  • Low Overlapping Scenes: The scene in Fig. 17(a) exhibits very low overlap in their subvolumes, resulting in a significant number of outliers between the non-overlapping regions. Due to the low inlier ratio, SG-Reg fails to successfully register the scenes.
  • Empty Room: SG-Reg relies on diverse semantic objects to build descriptive node features. The scene in Fig. 17(b) is an empty room devoid of furniture, containing only a few walls and a window with a noisy point cloud. Due to the lack of meaningful semantic objects, SG-Reg fails to identify any point inliers, resulting in registration failure.
  • Symmetric Subvolumes: The scene in Fig. 17(c) exhibits a symmetric room layout, with a table in the center and two identical couches on either side. Although SG-Reg finds correspondences with a 15.5% inlier ratio, many of these correspondences are adversarial, leading to a pseudo inlier ratio of $0\%$. Consequently, the registration fails.
To address the challenges shown in Fig. 17, there are a few possible direction in the future. 1) Improving semantic scene graph quality by constructing an implicit semantic map. Currently, we rely on FM-Fusion[11], which is an explicit semantic mapping. The ambiguous semantic categories, occlusion, and segmentation noise affect the reconstruction quality. In implicit semantic mapping, we can design a neural network to refine the semantic instance map in the 3D space. It should be able to further improve the reconstruction quality, which benefits the downstream registration task. 2) Extracting the appearance feature from each semantic node and fusing them into the node features. 3) Visual image registration, such as HLoc, does not conflict with scene graph registration. We may incorporate the image keyframes into our semantic scene graph. It jointly considers the image matching and semantic node matching to avoid those long-tailed cases.
Appendix E Two-agent SLAM
We supplement the implementation and additional results that have not been put in the main body due to page limits.
E-1 Compute HLoc bandwidth
We explain the HLoc configuration for calculating communication bandwidth in Table XII. In each keyframe, HLoc can extract fewer than $4096$ superpoints. Across all keyframe sequences, we have accumulated the total superpoints to be $460,000$. Each superpoint comprises a $256$-dimensional descriptor in float format, a $2$-dimensional coordinate in int format, and a scalar score in float. The total data size is calculated in megabytes (MB) and is presented in Table XII.
E-2 Discuss registration
NR(%) NP(%) IR(%) RR(%) RMSE(m)
Coarse Easy $83.9$ $88.4$ $83.5$ $0.08$
Median $68.0$ $68.7$ $51.2$ $0.11$
Difficult $42.3$ $26.2$ $0.0$
Dense Easy $86.1$ $89.0$ $6.4$ $100.0$ $0.05$
Median $77.5$ $75.5$ $5.8$ $75.0$ $0.08$
Difficult $62.1$ $35.3$ $1.4$ $0.0$
TABLE XVIII: Evaluate registration between the agents.
We also report our results following the registration metrics. We calculate the IoU of the input scene graphs at each query frame. The query frames are grouped into three sets: Difficult set has $\text{IoU}\in[0.1,0.3)$; Median set has $\text{IoU}\in[0.3,0.7)$; Easy set has $\text{IoU}\in[0.7,1.0)$. The RMSE threshold is $0.2$m. As shown in Table. XVIII, with coarse messages, our method has already registered $83.5\%$ of the easy scenes. Since we use a conservative registration strategy, our success rate is still low in the difficult set. The result further implies that the best way to apply SG-Reg in SLAM is to run registration at frames with large enough IoU. The failed registration at low IoU can be rejected by setting a threshold of the matched nodes.